The pressure in the atmosphere is not a constant but fluctuates as the weather changes. If the pressure outside a window drops by 10.1%, while the pressure inside does not change, what is the magnitude of the force on the window? Assume that the area of the window is 1.3 m2.
Answer #1

The magnetidue of the force on the window is

F = pA = 1.013 * 10^5 pa ( 0.101) ( 1.3 m^2 ) = 1.33 * 10 ^4 N